OpenStudy (anonymous):

can someone help me please (5/6 - 1/5) divided by 4/5

OpenStudy (aaronandyson):

\[\frac{ 5 }{ 6 } - \frac{ 1 }{ 5 }\]

OpenStudy (aaronandyson):

take LCM

OpenStudy (aaronandyson):

LCM = 30

OpenStudy (aaronandyson):

5 times 5 - 6 times 1 over 30

OpenStudy (aaronandyson):

25 - 6 / 30

OpenStudy (aaronandyson):

19/30

OpenStudy (aaronandyson):

so now 19/30 divided by 4/5

OpenStudy (anonymous):

wow your so awesome lol :) so do we multiply 19/30 by 4/5 now to get the answer

OpenStudy (aaronandyson):

Yes.

OpenStudy (anonymous):

okay so \[\frac{ 19 }{ 30 } \times \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

95/120 thats what i got i have a feeling i did something wrong tho lol

OpenStudy (aaronandyson):

Yep try again.

OpenStudy (anonymous):

76/150 now i got this

OpenStudy (anonymous):

@AaronAndyson

OpenStudy (anonymous):

is that wrong too ?

OpenStudy (aaronandyson):

Did you get 19/30 for the first part?

OpenStudy (aaronandyson):

5/6 take away 1/5

OpenStudy (anonymous):

no i didn't how did u get that lol ?

OpenStudy (aaronandyson):

I had explained it scroll up

OpenStudy (aaronandyson):

19/24 is the ans btw.

OpenStudy (anonymous):

ohhh ok thanks